OpenStudy (deadshot):

What is (f – g)(x) if f(x) = 2x – 3 and g(x) = 4x + 8?

OpenStudy (amistre64):

just subtract the functions

OpenStudy (deadshot):

so then it would be \[(2x-3)-(4x+8)=-2x-11\] right?

OpenStudy (amistre64):

f(x) = 2x – 3 g(x) = 4x + 8 -g(x) = -4x - 8 f(x) = 2x – 3 -g(x) = -4x - 8 --------------- (f-g)(x) = ??

OpenStudy (amistre64):

-2x - 11 is good

OpenStudy (deadshot):

Thanks!

OpenStudy (amistre64):

yw